OpenStudy (anonymous):

Please help: Limit of (x^3-4x)/(x^2-4) as x=2?

OpenStudy (johnweldon1993):

Have you studied L'Hopital's rule yet?

OpenStudy (anonymous):

factor and cacel

OpenStudy (anonymous):

Not really

OpenStudy (anonymous):

\[\frac{ x^3 - 4x }{ x^2 - 4 } = \frac{ x(x^2 - 4) }{ x^2 - 4 } = x\]

OpenStudy (anonymous):

ooop maybe "cancel"

OpenStudy (johnweldon1993):

Alright then follow @satellite73 and @Euler271 ...factor and cancel out

OpenStudy (anonymous):

Thank you